# TabSketchFM: Sketch-based Tabular Representation Learning for Data Discovery over Data Lakes

### v1

@@ 6.3. Application to search
Figures 11, 12, and 13, show that the fine tuned cross encoder models generalize rather well across tasks as well as domain, which is encouraging particularly because it shows that the models can be trained on a datalake that is quite different from the datalake that it is deployed to. In enterprise contexts, this is an important requirement, because it means that we can basically train models offline and apply it online. All models shown in the Figure are models that include the value embeddings from a column for maximal generalization, but using the models without values also produced the same type of generalization.
Figure 11. F1 on wiki join transfer on search for varying k.
Figure 12. F1 on Santos transfer on search for varying k.
Figure 13. F1 on TUS transfer on search for varying k.
7. Conclusion and Discussion
We presented TabSketchFM, a novel transformer-based tabular representation model that inputs the table in the form of several sketches. We empirically show that cross encoders based on TabSketchFM outperform tabular models that serialize tables as texts, for three different data discovery tasks. We also show that the finetuned TabSketchFM can be effectively used to rerank the union and join search results. To further the research, we share 8 new datasets (LakeBench) that can be used to finetune neural models to understand tables for data discovery tasks. In the future, we will identify other data discovery tasks that TabSketchFM can be useful, and also make the model more robust for example, by incorporating other hashing mechanisms and cell values to create the sketches.

### v3

@@ IV-C4 Generalization across tasks and domains
We now investigate Q4. How does our method, fine-tuned on a single task and domain, can adapt and generalize to a new domain. Figures 8(a), 8(b), and 8(c), 8(d), show that the fine tuned cross encoder models generalize rather well across tasks as well as domain, which is encouraging particularly because it shows that the models can be trained on a datalake that is quite different from the datalake that it is deployed to. In enterprise contexts, this is an important requirement, because it means that we can basically train models offline and apply it online. All models shown in the Figure are models that include the value embeddings from a column for maximal generalization, but using the models without values also produced the same type of generalization.
V Conclusion and Discussion
We presented TabSketchFM, a novel transformer-based tabular representation model that inputs the table in the form of several sketches. We show that cross encoders based on TabSketchFM outperform existing systems on LakeBench related table identification task. We also show that finetuned TabSketchFM can be effectively used to search for tables that are unionable, joinable or are subsets of each other. We highlight that embedding values using sentence encoder models can be a useful addition for some discovery tasks. As far as the search efficiency is concerned, we recommend indexing the datalake offline and at query time only compute embeddings for the query table. This computation is similar to other embedding-based search system.